An electron is moving in the 3rd orbit of hydrogen atom and radius of first orbit is x. Then.
  1. de-broglie wavelength is 6pix
  2. de-broglie wavelength is 2pix
  3. electron velocity is h/6pixm
  4. electron velocity is h/2pixm

Grade:11

1 Answers

Sakshi
8 years ago
Given that the radius of first orbit is x,
one can use the formula
r=(0.529*n2)/z
putting n=z=1
we can calculate r and that value of r will be x
now, since the electron is moving in 3rd orbit,
one can cal r from the above formula again with n=3
to get the ans try putting the value of r in the formula for velocity, as
v=(n*h)/(2*3.14*m*r)
or
further you can find wavelength = (h/m*v)